OpenStudy (anonymous):

find each unit cost $12 for 4 yards to the 2 power $3.45 for 3.7 oz $9 for 5 L

OpenStudy (saifoo.khan):

12/4 = 3 3.45/3.7 = 0.93 9/5 = 1.8

OpenStudy (anonymous):

O.o tofu!

OpenStudy (anonymous):

WAIT What -.-

OpenStudy (anonymous):

im confused on the fractio one XD

OpenStudy (saifoo.khan):

all three are fractions ones. lol

OpenStudy (anonymous):

OH LOL!

OpenStudy (anonymous):



OpenStudy (saifoo.khan):

you want the cost for 1 unit. so in the first one, u have the cost for 4 units. so for 1 unit will be: 12/4 = 3yards^2

OpenStudy (anonymous):

TOFU!!!! new name for you i guess. cute

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

so 3 yards

OpenStudy (saifoo.khan):

yes!!

OpenStudy (saifoo.khan):

and same for others

OpenStudy (anonymous):

SO 3 FOR ALL =D

OpenStudy (saifoo.khan):

No, 3 for the first one.

OpenStudy (saifoo.khan):

3.45/3.7 = 0.93 9/5 = 1.8